OpenStudy (anonymous):

Find a quadratic equation in standard form in standard form with integral coefficients and having the give numbers as its roots.

OpenStudy (anonymous):

\[3-2\sqrt{7} , 3+2\sqrt{7}\]

OpenStudy (saifoo.khan):

x = 3 - 2sqrt7 and x = 3 + 2 sqrt 7\ x -(3-2sqrt7) and x -(3+ 2sqrt7) Now multiply them both: \[[x-(3-2\sqrt7)] \times [x - (3+2\sqrt7)]\]

OpenStudy (saifoo.khan):

You should get: \[\Large x^2-6x-19\]

OpenStudy (anonymous):

ahhh. Thank you.

OpenStudy (saifoo.khan):

You're welcome.